OpenStudy (anonymous):

Two electric bulbs whose resistances are in the ratio of 1 : 2 are connected in parallel to a constant voltage source. The powers dissipated in them have the ratio. a 1 : 2 b 1 : 1 c 2 : 1 d 1 : 4

OpenStudy (anonymous):

Using I*V=p we know that in a parallel circuit the voltage is always the same but the current is different if we use I=V/R and take for example a 4ohm resistor and and 8 ohm resistor, lets say the voltage is 3 volts I=3/4 I=.75 amps I=3/8 I=.375amps So if we plug it back into the equation I*V=p (.375)*3=1.125 (.75)*3=2.25 you can now see that it is a 2:1 relationship